# Bogomolov property and Galois representations

### v1

@@ 1. Introduction
We fix two (potentially infinite) Galois extensions $F/{\mathbb{Q}}$, $L/{\mathbb{Q}}$ with $F\subseteq L$. We fix a rational prime $p$. Usually lower bound for the height comes from a dicotomy depending on whether $p$ is ramified or not in $L$. In the second case a Frobenius argument is generally enough to conclude. In the first case we need a descent argument. Here we change the point of view, assuming $F$ not ramified at $p$, with the residual degrees at $p$ not growing from $F$ to $L$. Moreover, we avoid the equidistribution argument of [15], by using instead the acceleration lemma of [1]. This simplifies the proofs, although it leads to worse lower bounds, as remarked in [12, Section 4.2.2].
We start with a simple result, which a special case of  [1, Theorem 1.5]. Here we allow finite ramification at $p$ in $F$.
Proposition 2.1.
Let $G=\mathrm{Gal}(F/{\mathbb{Q}})$ and let $Z$ be the center. We assume that the ramification index of $p$ in $F$ and the order of a lift of a Frobenius $\sigma$ over $p$ in $G/Z$ are both finite. Then $F$ has property (B).
Proof.
Let $e$ and $r$ be respectively the ramification index and the order of $\sigma$ in $G/Z$. Let $F_{0}/{\mathbb{Q}}$ be a finite Galois extension and $\alpha\in F_{0}$. Since $\sigma^{r}$ lies in the center of $\mathrm{Gal}(F_{0}/{\mathbb{Q}})$, we have
$|\sigma^{r}(\gamma)-\gamma^{p^{r}}|_{v}\leq p^{-1/e}$
for any $\gamma\in{\mathcal{O}}_{F_{0}}$ and for any place $v$ dividing $p$. By [1, Lemma 2.1], there exists a positive integer $\lambda$ which is explicitly bounded in terms of $p$ and of $e$ such that
$|\sigma^{r}(\gamma^{p^{\lambda}})-\gamma^{p^{r+\lambda}}|_{v}\leq p^{-1}$
We now make the first assumption on $L$ which allow to globalise the metric inequality at ramified primes. We follow some ideas from [3, Lemma 4.2] and [15, Proof of Proposition 6.1]. Let $\epsilon_{p}:G_{\mathbb{Q}}\to\mathbb{Z}_{p}^{\times}$ denote the $p$-adic cyclotomic character.
Assumption 2.2.
There exists $\tau\in G_{\mathbb{Q}}\cap I$, where $I$ is an inertia group at $p$, such that
(2.1) $\epsilon_{p}(\tau)=g\quad\hbox{with }g\in\mathbb{Z},g>1$
and $\tau|_{L}\in Z(\mathrm{Gal}(L/{\mathbb{Q}}))$.
The second assumption concerns ramification groups. Although of technical nature, it is quite general to be satisfied in most cases, namely for the extensions cut out by our Galois representations. Even we are not interested in explicit computations, let’s remark that some of these assumptions can be weakened in some cases to get sharper lower bounds. Given a finite Galois extension $E/{\mathbb{Q}}$ and $i\geq 0$ we recall that the $i$-th ramification group $G_{i}$ at a place $v$ over $p$ is the subgroup of $\sigma\in\mathrm{Gal}(E/{\mathbb{Q}})$ such that $|\sigma(\gamma)-\gamma|_{v}\leq p^{-(i+1)/e}$ for $\gamma\in{\mathcal{O}}_{E}$. Here $e$ is the ramification index of $E$ over $p$. The group $G_{i}$ is the isomorphic image, by restriction, of the corresponding local $i$-th ramification group of the completed extension $E_{v}/\mathbb{Q}_{p}$. As it is well known, the $i$-th ramification groups at different places of $E$ over $p$ are conjugate.
Assumption 2.3.
There exist $C_{1}$, $C_{2}\geq 1$ with the following properties. There exists a sequence of finite Galois extensions $\{L_{n}/{\mathbb{Q}}\}_{n\geq 0}$ with
${\mathbb{Q}}=L_{0}\subsetneq L_{1}\subsetneq\cdots\subsetneq L_{n}\subsetneq\cdots\subsetneq L,\quad{\rm and}\quad L=\cup_{n}L_{n}.$
such that for $n\geq 1$ the extension $L_{n}/{\mathbb{Q}}$ is ramified over $p$. Moreover, let $e_{n}\geq 1$ be the ramification index, $(G^{n}_{i})_{i}$ the sequence of ramification groups of $L_{n}/{\mathbb{Q}}$ at a place over $p$. We further denote by $G^{n}_{i_{n}}$ the last non trivial ramification group and $H_{n}$ its normal closure in $\mathrm{Gal}(L_{n}/{\mathbb{Q}})$. Then
$\frac{e_{n}}{i_{n}+1}\leq C_{1},\qquad|H_{n}|\leq C_{2}$
and
$L_{n}^{H_{n}}=L_{n-1}.$
We can now state the main result of this section. We denote by ${\boldsymbol{\mu}}$ the set of roots of unity and by ${\boldsymbol{\mu}}_{p^{\infty}}$ the set of roots of unity of order a power of $p$. For the reader convenience, we also recall a statement from [15, Lemma 2.1].
@@
Let us assume 2.2 and 2.3. If $F$ is unramified at $p$ and has property (B), then $FL$ has property (B) as well.
@@
Let $I\subseteq\mathrm{Gal}(L/{\mathbb{Q}})$ be an inertia group at $p$, $\tau\in I$ and $g\in{\mathbb{Z}}$ as in Assumption 2.2. Remark that $\tau$ is trivial on $F$, since $\tau$ is in an inertia group and $F$ is not ramified at $p$. Since $F$ has property (B), there exists $c_{0}>0$ such that
(2.2) $\forall\beta\in F^{\star}\backslash{\boldsymbol{\mu}},\quad h(\beta)\geq c_{0}.$
Let $\alpha\in(FL)^{\star}\backslash{\boldsymbol{\mu}}$. We define
@@
We fix a finite Galois extension $F_{0}/{\mathbb{Q}}$ such that $F_{0}\subseteq F$ and $\beta\in F_{0}L$. Let $\{L_{n}/{\mathbb{Q}}\}_{n\geq 0}$ such that Assumption 2.3 is satisfied. We define $n\geq 0$ as the smallest integer such that $\beta\in F_{0}L_{n}$. If $n=0$ then $\beta\in F$ and $h(\beta)\geq c_{0}$ by (2.2), since $\beta\not\in{\boldsymbol{\mu}}$. Thus we can assume $n\geq 1$.
We now fix a place $v$ of $F_{0}L_{n}$ over $p$. Since $F_{0}$ is not ramified at $p$, the ramification index $e_{n}$ of $v_{|L_{n}}$ coincide with the ramification index of $v$. Let $G^{n}_{i}\subseteq\mathrm{Gal}(L_{n}/{\mathbb{Q}})$ and $G^{\prime n}_{i}\subseteq\mathrm{Gal}(F_{0}L_{n}/{\mathbb{Q}})$ be respectively the $i$-th ramification subgroups at $v_{|L_{n}}$ and at $v$. The restriction map
$\mathrm{Gal}(F_{0}L_{n}/F_{0}L_{n-1})\rightarrow\mathrm{Gal}(L_{n}/L_{n-1})$
is injective; by Lemma 2.4 it sends $G^{\prime n}_{i_{n}}$ bijectively in $G^{n}_{i_{n}}$. Therefore the normal closure $H^{\prime}_{n}$ of $G^{\prime n}_{i_{n}}$ in $\mathrm{Gal}(F_{0}L_{n}/{\mathbb{Q}})$ is sent bijectively in the normal closure $H_{n}$ of $G^{n}_{i_{n}}$ in $\mathrm{Gal}(L_{n}/{\mathbb{Q}})$. By Assumption 2.3 we have
(2.4) $\frac{e_{n}}{i_{n}+1}\leq C_{1},\qquad|H^{\prime}_{n}|\leq C_{2}.$
Since $L_{n}^{H_{n}}=L_{n-1}$ by the last claim of Assumption 2.3, Galois theory shows that
(2.5) $(F_{0}L_{n})^{H^{\prime}_{n}}=F_{0}L_{n-1}.$
Thus $\beta\not\in(F_{0}L_{n})^{H^{\prime}_{n}}$ by minimality of $n$. Since $H^{\prime}_{n}$ is generated by the conjugates of $G^{\prime n}_{i_{n}}$, by Galois theory $\beta\not\in(F_{0}L_{n})^{\phi G^{\prime n}_{i_{n}}\phi^{-1}}$ for some $\phi\in\mathrm{Gal}(F_{0}L_{n}/{\mathbb{Q}})$. Thus there exists $\sigma\in G^{\prime n}_{i_{n}}$ such that $(\phi\sigma\phi^{-1})(\beta)\neq\beta$. We fix $\phi$, $\sigma$ with the above property and we let $\beta^{\prime}=\phi^{-1}(\beta)$. Thus $\sigma(\beta^{\prime})\neq\beta^{\prime}$. Moreover, if $\tilde{\phi}$ is a lifting of $\phi$ to $FL$, we have $\beta^{\prime}=\tfrac{\tau(\alpha^{\prime})}{\alpha^{\prime g}}$ with $\alpha^{\prime}=\tilde{\phi}^{-1}(\alpha)$, by definition (2.3) and since $\tau$ is in the center of $\mathrm{Gal}(FL/{\mathbb{Q}})$. Note also that $\beta^{\prime}\in F_{0}L_{n}$ and $h(\beta^{\prime})=h(\beta)$ since $\beta^{\prime}$ is a conjugate of $\beta$. Summing up, we have found $\sigma\in G^{\prime n}_{i_{n}}$ and $\alpha^{\prime}\in FL$ such that
@@
Let us assume by contradiction that $\tfrac{\sigma(\beta^{\prime})}{\beta^{\prime}}\in{\boldsymbol{\mu}}_{p^{\infty}}$. Let $\tilde{\sigma}$ be a lift of $\sigma$ to $FL$. By (2.6) and since $\tau$ is central,
@@
with $\eta=\tilde{\sigma}(\alpha^{\prime})/\alpha^{\prime}\in FL$. Hence
@@
Since $\sigma\in G^{\prime n}_{i_{n}}$, for any $\gamma\in{\mathcal{O}}_{F_{0}L_{n}}$ we have
@@
(2.8) $|\sigma(\gamma^{p^{\lambda}})-\gamma^{p^{\lambda}}|_{v}\leq p^{-|H^{\prime}_{n}|}.$
We consider the action of $\mathrm{Gal}(F_{0}L_{n}/{\mathbb{Q}})$ on itself by conjugation and its stabiliser $C(\sigma)=\{\tau\in\mathrm{Gal}(F_{0}L_{n}/{\mathbb{Q}})\;|\;\tau\sigma\tau^{-1}=\sigma\}$. Let $\tau\in C(\sigma)$. We use (2.8) with $\tau^{-1}\gamma$ instead of $\gamma$. Since $\sigma\tau^{-1}=\tau^{-1}\sigma$ and $|\tau^{-1}(\star)|_{w}=|\star|_{\tau(v)}$, we get
$|\sigma(\gamma^{p^{\lambda}})-\gamma^{p^{\lambda}}|_{w}\leq p^{-|H^{\prime}_{n}|}$
with $w=\tau(v)$ and thus for any $w$ in the orbit $S$ of $v$ under the the action of $C(\sigma)$ on the places avove $p$. Using [2, Lemma 1] as in the proof of [1, Lemma 2.1] we find
@@
with $c(w)=p^{-|H^{\prime}_{n}|}$. This last inequality also holds for an arbitrary place $w$ of $L$ with
@@
Let $d_{n}=[F_{0}L_{n}:{\mathbb{Q}}]$ and $d_{n,w}=[(F_{0}L_{n})_{w}:{\mathbb{Q}}_{w}]$ with $w$ a place of $F_{0}L_{n}$. Using the Product Formula to $\sigma(\beta^{\prime})^{p^{\lambda}}-\beta^{\prime p^{\lambda}}$, which is $\neq 0$ by (2.7), as in the proof of [1, Lemma 2.1] we get:

### v3

@@ 1. Introduction
Section 3 is devoted to two general statements on lower bounds for the height. The proofs rely on the techniques developed in [2], [3], [15] and [1]. We fix two (possibly infinite) Galois extensions ${\mathbb{F}}/{\mathbb{Q}}$, ${\mathbb{L}}/{\mathbb{Q}}$. We fix a rational prime $p$. Usually, lower bounds for the height come from a dicotomy depending on whether $p$ is ramified or not in ${\mathbb{L}}$. In the second case a Frobenius argument is generally enough to conclude. In the first case we need a descent argument. Here we change the point of view, assuming ${\mathbb{F}}$ not ramified at $p$, with the residual degrees at $p$ not growing from ${\mathbb{F}}$ to ${\mathbb{L}}$. Moreover, we avoid the equidistribution argument of [15], by using instead the acceleration lemma of [1]. This simplifies the proofs, although it leads to worse lower bounds, as remarked in [12, Section 4.2.2]. As a toy example we deduce a generalization (Theorem 3.5) of the main result of [2].
One of the crucial steps in applying the height-machinery above is to have a precise as possible description of the sequence of the local higher ramification groups. This is done in [15, Lemma 3.3] using the fact that the formal group associated to the elliptic curve is a Lubin-Tate module over ${\mathbb{Z}}_{p}$. In Section 4 we show that, by local Langlands correspondence and the classification of the two dimensional crystalline representations of $G_{\mathbb{Q}_{p}}$ given by Breuil, our local representation at $p$ is a twist of a power of a fundamental character, so that Lubin Tate theory is also applicable in our more general case.
Section 5 is devoted to prove Proposition 5.1, which is the analogous of Lemma 6.2 of [15] on the normal closure of the last ramification group in the field cut out by the reduction modulo $p^{n}$ of our modular representation. This was possible thanks to condition $(P2)$ in Assumption 1.3. For the proof we needed a group theoretical study of the linear group of a finite product of local rings. This is done in Appendix A.
The concluding Section 6 contains the proof of the results announced in the introduction.
2. Notations and preliminary results
In the following
  1. -
    local fields will be denoted by $F,K,E,L,\ldots$
  2. -
    global fields will be denoted by $\mathbb{F},\mathbb{K},\mathbb{E},\mathbb{L},\ldots$
For a field $k$, $\overline{k}$ denotes an algebraic closure of $k$, and $G_{k}=\mathrm{Gal}(\overline{k}/k)$ the absolute Galois group of $k$, endowed with the profinite topology.
Let $\Omega$ be a separated topological group and $\rho:G_{\mathbb{F}}\to\Omega$ be a continuous homomorphism; we shall denote by $k(\rho)=\overline{k}^{\ker\rho}$ the field cut out by $\rho$. Then $k(\rho)/k$ is a Galois extension whose Galois group is isomorphic to ${\rm Im}(\rho)$. Moreover, since $\ker{\rho}$ is closed, $G_{k(\rho)}=\ker\rho$.
When $k$ is a local or global field and $v$ is a place of $k$, we shall say that $\rho$ is unramified (resp. totally ramified) at $v$ if $k(\rho)$ is. Note that $\rho$ induces an isomorphism $\mathrm{Gal}(k(\rho)/k)\rightarrow{\rm Im}(\rho)$ which as usual we denote by the same letter $\rho$.
From now on we fix a normalized new eigenform $f=\sum_{n}a_{n}q^{n}\in S_{k}(\Gamma_{0}(N))$ for the Hecke algebra of weight $k\geq 2$; then $k$ is an even integer. We recall the relevant definitions. We denote by ${\mathbb{K}}$ the number field generated by the $a_{n}$’s and by $\mathcal{O}$ its integer ring. A construction due to Shimura and Deligne associates to $f$ a family of compatible Galois representations:
$\rho_{v}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\mathcal{O}_{v}),\quad v\hbox{ finite place of }{\mathbb{K}}.$
[…]
(2.2) for Npv, the characteristic polynomial of the image by ρv ofa Frobenius element at  is $\hbox{for $\ell\nmid Np_{v}$, the characteristic polynomial of the image by $\rho_{v}$ of}\\ \quad\quad\quad\hbox{a Frobenius element at $\ell$ is $X^{2}-a_{\ell}X+\ell^{k-1}$.}\hfill$
Remark 2.1.
Let $p$ be a prime and $v$ be any place of ${\mathbb{K}}$ dividing $p$. Then, by (2.2) and by Chebotarev lemma, the determinant of $\rho_{v}$ is ${\varepsilon}_{p}^{k-1}$, where ${\varepsilon}_{p}\colon G_{\mathbb{Q}}\rightarrow{\mathbb{Z}}_{p}^{\times}$ is the $p$-adic cyclotomic character. (Indeed, for a Frobenius element $\mathrm{Frob}_{\ell}$ with $\ell\nmid Np$, we have: $\det\big{(}\rho_{v}(\mathrm{Frob}_{\ell})\big{)}=\ell^{k-1}={\varepsilon}_{p}^{k-1}(\mathrm{Frob}_{\ell})$.)
We let
$\rho=\prod_{v}\rho_{v}:G_{\mathbb{Q}}\longrightarrow\mathrm{GL}_{2}(\widehat{\mathcal{O}}),$
where $\widehat{\mathcal{O}}$ is the profinite completion of $\mathcal{O}$. For a prime $p$ and $n>1$, we also denote by $\rho(p^{n})$ the reduction $G_{\mathbb{Q}}\to\mathrm{GL}_{2}(\widehat{\mathcal{O}}/p^{n}\widehat{\mathcal{O}})=\mathrm{GL}_{2}(\mathcal{O}/p^{n}\mathcal{O})$ of $\rho$.
Remark 2.2.
Let $p$ be a prime which does not divide $N$ and assume that $p=\prod\limits_{v\mid p}{\mathfrak{P}}_{v}^{e_{v}}$ in $\mathcal{O}$. Let $n\geq 1$ be an integer. Then
                  $G_{\mathbb{Q}}$$\prod\limits_{v}\mathrm{GL}_{2}(\mathcal{O}_{v})$$\prod\limits_{v\mid p}\mathrm{GL}_{2}(\mathcal{O}_{v}/{\mathfrak{P}}_{v}^{e_{v}n})$,$\mathrm{GL}_{2}(\widehat{\mathcal{O}})$$\mathrm{GL}_{2}(\mathcal{O}/p^{n}\mathcal{O})$$\rho=\prod\limits_{v}\rho_{v}$$\pi$$\rho(p^{n})$
[…]
We let
For the reader convenience, we also recall a statement from [15, Lemma 2.1].
@@ 2. Notations and preliminary results
3. Diophantine results.
We start with a simple result, which is a special case of  [1, Theorem 1.5]. Here we allow finite ramification at $p$ in ${\mathbb{F}}$.
@@
Let $G=\mathrm{Gal}({\mathbb{F}}/{\mathbb{Q}})$ and let $Z$ be the center. We assume that the ramification index of $p$ in ${\mathbb{F}}$ and the order of a lift of a Frobenius $\sigma$ over $p$ in $G/Z$ are both finite. Then ${\mathbb{F}}$ has property (B).
@@
Let $e$ and $r$ be respectively the the ramification index and the order of $\sigma$ in $G/Z$. Let ${\mathbb{F}}_{0}/{\mathbb{Q}}$ be a finite Galois extension and $\alpha\in{\mathbb{F}}_{0}$. Since $\sigma^{r}$ lies in the center of $\mathrm{Gal}({\mathbb{F}}_{0}/{\mathbb{Q}})$, we have
$|\sigma^{r}(\beta)-\beta^{p^{r}}|_{v}\leq p^{-1/e}$
for any $\beta\in{\mathcal{O}}_{{\mathbb{F}}_{0}}$ and for any place $v$ dividing $p$. By [1, Lemma 2.1], there exists a positive integer $\lambda$ which is explicitly bounded in terms of $p$ and of $e$ such that
$|\sigma^{r}(\beta^{p^{\lambda}})-\beta^{p^{r+\lambda}}|_{v}\leq p^{-1}$
Let ${\mathbb{L}}$ be a (possible infinite) Galois extension. We now make the first assumption on ${\mathbb{L}}$ which allows to globalise the metric inequality at ramified primes. We follow some ideas from [3, Lemma 4.2] and [15, Proof of Proposition 6.1].
Assumption 3.2.
There exists $\tau\in G_{\mathbb{Q}}\cap I$, where $I$ is an inertia group [^3] That is, the inverse limit of inertia groups. over $p$, such that
(3.1) $\epsilon_{p}(\tau)=g\quad\hbox{with }g\in\mathbb{Z},g>1,$
and $\tau|_{\mathbb{L}}\in Z(\mathrm{Gal}({\mathbb{L}}/{\mathbb{Q}}))$.
The second assumption on ${\mathbb{L}}$ concerns ramification groups. Although of technical nature, it is quite general to be satisfied in several cases, namely for the extensions cut out by our Galois representations. Given a finite Galois extension ${\mathbb{E}}/{\mathbb{Q}}$ and $i\geq 0$ we recall that the $i$-th ramification group $G_{i}$ at a place $v$ over $p$ is the subgroup of $\sigma\in\mathrm{Gal}({\mathbb{E}}/{\mathbb{Q}})$ such that $|\sigma(\gamma)-\gamma|_{v}\leq p^{-(i+1)/e}$ for $\gamma\in{\mathcal{O}}_{\mathbb{E}}$. Here $e$ is the ramification index of ${\mathbb{E}}$ over $p$. The group $G_{i}$ is the isomorphic image, by restriction, of the corresponding local $i$-th ramification group of the completed extension ${\mathbb{E}}_{v}/\mathbb{Q}_{p}$. As it is well known, the $i$-th ramification groups at different places of ${\mathbb{E}}$ over $p$ are conjugate.
Assumption 3.3.
There exist $C_{1}$, $C_{2}\geq 1$ with the following properties. There exists a sequence of finite Galois extensions $({\mathbb{L}}_{n}/{\mathbb{Q}})_{n\geq 0}$ with
${\mathbb{Q}}={\mathbb{L}}_{0}\subsetneq{\mathbb{L}}_{1}\subsetneq\cdots\subsetneq{\mathbb{L}}_{n}\subsetneq\cdots\subsetneq{\mathbb{L}},\quad{\rm and}\quad{\mathbb{L}}=\bigcup_{n}{\mathbb{L}}_{n}.$
such that for $n\geq 1$ the extension ${\mathbb{L}}_{n}/{\mathbb{Q}}$ is ramified over $p$. Moreover, let $e_{n}\geq 1$ be the ramification index, $(G^{n}_{i})_{i}$ the sequence of ramification groups of ${\mathbb{L}}_{n}/{\mathbb{Q}}$ at a place over $p$. We further denote by $H_{n}=G^{n}_{i_{n}}$ the last non trivial ramification group and $\overline{H_{n}}$ its normal closure in $\mathrm{Gal}({\mathbb{L}}_{n}/{\mathbb{Q}})$. Then
$\frac{e_{n}}{i_{n}+1}\leq C_{1},\qquad|\overline{H_{n}}|\leq C_{2}$
and
${\mathbb{L}}_{n}^{\overline{H_{n}}}={\mathbb{L}}_{n-1}.$
Even we are not interested in explicit computations, let us remark that the above assumptions can be weakened in some cases (as for abelian extensions) to get sharper lower bounds.
We can now state the main result of this section. We denote by ${\boldsymbol{\mu}}$ the set of roots of unity and by ${\boldsymbol{\mu}}_{p^{\infty}}$ the set of roots of unity of order a power of $p$ in $\overline{\mathbb{Q}}$.
Proposition 3.4.
Let ${\mathbb{F}}/{\mathbb{Q}}$ and ${\mathbb{L}}/{\mathbb{Q}}$ two (possibly infinite) Galois extensions. We suppose that:
  1. -
    ${\mathbb{F}}$ is unramified at $p$ and it satisfies property (B);
  2. -
    ${\mathbb{L}}$ satisfies Assumptions  3.2 and 3.3.
Then ${\mathbb{F}}{\mathbb{L}}$ has property (B) as well.
Proof.
Let $I\subseteq\mathrm{Gal}(\overline{\mathbb{Q}}/{\mathbb{Q}})$ be an inertia group over $p$, $\tau\in I$ and $g\in{\mathbb{Z}}$ as in Assumption 3.2.
By definition of inertia group and by Assumption 3.2, there exists a prime ideal ${\mathfrak{P}}$ of ${\mathbb{F}}$ over $p$ such that $\tau_{|{\mathbb{F}}}\in I({\mathfrak{P}}|p)$. Since ${\mathbb{F}}$ is not ramified at $p$, this last group is trivial, thus $\tau$ is trivial on ${\mathbb{F}}$. Again by Assumption 3.2, $\tau|_{\mathbb{L}}\in Z(\mathrm{Gal}({\mathbb{L}}/{\mathbb{Q}}))$. Thus
(3.2) $\tau|_{{\mathbb{F}}{\mathbb{L}}}\in Z(\mathrm{Gal}({\mathbb{F}}{\mathbb{L}}/{\mathbb{Q}})).$
Since ${\mathbb{F}}$ has property (B), there exists $c_{0}>0$ such that
(3.3) $\forall\beta\in{\mathbb{F}}^{\star}\backslash{\boldsymbol{\mu}},\quad h(\beta)\geq c_{0}.$
Let $\alpha\in({\mathbb{F}}{\mathbb{L}})^{\star}\backslash{\boldsymbol{\mu}}$. We define
@@
We fix a finite Galois extension ${\mathbb{F}}_{0}/{\mathbb{Q}}$ such that ${\mathbb{F}}_{0}\subseteq{\mathbb{F}}$ and $\beta\in{\mathbb{F}}_{0}{\mathbb{L}}$. Let $\{{\mathbb{L}}_{n}/{\mathbb{Q}}\}_{n\geq 0}$ as in Assumption 3.3. We define $n\geq 0$ as the smallest integer such that $\beta\in{\mathbb{F}}_{0}{\mathbb{L}}_{n}$. If $n=0$ then $\beta\in{\mathbb{F}}$ and $h(\beta)\geq c_{0}$ by (3.3), since $\beta\not\in{\boldsymbol{\mu}}$. Thus we can assume $n\geq 1$.
We now fix a place $v$ of ${\mathbb{F}}_{0}{\mathbb{L}}_{n}$ over $p$. Since ${\mathbb{F}}_{0}$ is not ramified at $p$, the ramification index $e_{n}$ of $v_{|{\mathbb{L}}_{n}}$ coincide with the ramification index of $v$. Let $G^{n}_{i}\subseteq\mathrm{Gal}({\mathbb{L}}_{n}/{\mathbb{Q}})$ and $G^{\prime n}_{i}\subseteq\mathrm{Gal}({\mathbb{F}}_{0}{\mathbb{L}}_{n}/{\mathbb{Q}})$ be respectively the $i$-th ramification subgroups at $v_{|{\mathbb{L}}_{n}}$ and at $v$. The restriction map $\mathrm{Gal}({\mathbb{F}}_{0}{\mathbb{L}}_{n}/{\mathbb{F}}_{0})\rightarrow\mathrm{Gal}({\mathbb{L}}_{n}/\mathbb{Q})$ sends $H^{\prime}_{n}:=G^{\prime n}_{i_{n}}$ bijectively in $H_{n}=G^{n}_{i_{n}}$, by Lemma 2.3 applied to the completions at $v$ of the involved fields. Therefore the normal closure $\overline{H^{\prime}_{n}}$ of $H^{\prime}_{n}$ in $\mathrm{Gal}({\mathbb{F}}_{0}{\mathbb{L}}_{n}/{\mathbb{Q}})$ is sent bijectively in the normal closure $\overline{H_{n}}$ of $H_{n}$ in $\mathrm{Gal}({\mathbb{L}}_{n}/{\mathbb{Q}})$. By Assumption 3.3 we have
(3.5) $\frac{e_{n}}{i_{n}+1}\leq C_{1},\qquad|\overline{H^{\prime}_{n}}|\leq C_{2}.$
Since ${\mathbb{L}}_{n}^{\overline{H_{n}}}={\mathbb{L}}_{n-1}$ by the last claim of Assumption 3.3, Galois theory shows that
(3.6) $({\mathbb{F}}_{0}{\mathbb{L}}_{n})^{\overline{H^{\prime}_{n}}}={\mathbb{F}}_{0}{\mathbb{L}}_{n-1}.$
Thus $\beta\not\in({\mathbb{F}}_{0}{\mathbb{L}}_{n})^{\overline{H^{\prime}_{n}}}$ by minimality of $n$. Since $\overline{H^{\prime}_{n}}$ is generated by the conjugates of $G^{\prime n}_{i_{n}}$, by Galois theory $\beta\not\in({\mathbb{F}}_{0}{\mathbb{L}}_{n})^{\phi G^{\prime n}_{i_{n}}\phi^{-1}}$ for some $\phi\in\mathrm{Gal}({\mathbb{F}}_{0}{\mathbb{L}}_{n}/{\mathbb{Q}})$. Thus there exists $\sigma\in G^{\prime n}_{i_{n}}$ such that $\phi\sigma\phi^{-1}(\beta)\neq\beta$. We fix $\phi$, $\sigma$ with the above property and we let $\beta^{\prime}=\phi^{-1}(\beta)$. Thus $\sigma(\beta^{\prime})\neq\beta^{\prime}$. Moreover, if $\tilde{\phi}$ is a lifting of $\phi$ to ${\mathbb{F}}{\mathbb{L}}$, we have $\beta^{\prime}=\tfrac{\tau(\alpha^{\prime})}{\alpha^{\prime g}}$ with $\alpha^{\prime}=\tilde{\phi}^{-1}(\alpha)$, by definition (3.4) and since the restriction of $\tau$ is in the center of $\mathrm{Gal}({\mathbb{F}}{\mathbb{L}}/{\mathbb{Q}})$ (cf. (3.2)). Note also that $\beta^{\prime}\in{\mathbb{F}}_{0}{\mathbb{L}}_{n}$ and $h(\beta^{\prime})=h(\beta)$ since $\beta^{\prime}$ is a conjugate of $\beta$. Summing up, we have found $\sigma\in G^{\prime n}_{i_{n}}$ and $\alpha^{\prime}\in{\mathbb{F}}{\mathbb{L}}$ such that
@@
Let us assume by contradiction that $\tfrac{\sigma(\beta^{\prime})}{\beta^{\prime}}\in{\boldsymbol{\mu}}_{p^{\infty}}$. Let $\tilde{\sigma}$ be a lift of $\sigma$ to ${\mathbb{F}}{\mathbb{L}}$. By (3.7) and since $\tau$ is central,
@@
with $\eta=\tilde{\sigma}(\alpha^{\prime})/\alpha^{\prime}\in{\mathbb{F}}{\mathbb{L}}$. Hence
@@
Since $\sigma\in G^{\prime n}_{i_{n}}$, for any $\gamma\in{\mathcal{O}}_{{\mathbb{F}}_{0}{\mathbb{L}}_{n}}$ we have
@@
(3.9) $|\sigma(\gamma^{p^{\lambda}})-\gamma^{p^{\lambda}}|_{v}\leq p^{-|\overline{H^{\prime}_{n}}|}.$
We consider the action of $\mathrm{Gal}({\mathbb{F}}_{0}{\mathbb{L}}_{n}/{\mathbb{Q}})$ on itself by conjugation and its stabiliser $C(\sigma)=\{\tau\in\mathrm{Gal}({\mathbb{F}}_{0}{\mathbb{L}}_{n}/{\mathbb{Q}})\;|\;\tau\sigma\tau^{-1}=\sigma\}$. Let $\tau\in C(\sigma)$. We use (3.9) with $\tau^{-1}\gamma$ instead of $\gamma$. Since $\sigma\tau^{-1}=\tau^{-1}\sigma$ and $|\tau^{-1}(\star)|_{w}=|\star|_{\tau(v)}$, we get
$|\sigma(\gamma^{p^{\lambda}})-\gamma^{p^{\lambda}}|_{w}\leq p^{-|\overline{H^{\prime}_{n}}|}$
with $w=\tau(v)$ and thus for any $w$ in the orbit $S$ of $v$ under the the action of $C(\sigma)$ on the places above $p$. Using [2, Lemma 1] as in the proof of [1, Lemma 2.1] we find
@@
with $c(w)=p^{-|\overline{H^{\prime}_{n}}|}$. This last inequality also holds for an arbitrary place $w$ of ${\mathbb{L}}$ not dividing $p$, with
@@
Let $d_{n}=[{\mathbb{F}}_{0}{\mathbb{L}}_{n}:{\mathbb{Q}}]$ and $d_{n,w}=[({\mathbb{F}}_{0}{\mathbb{L}}_{n})_{w}:{\mathbb{Q}}_{w}]$ with $w$ a place of ${\mathbb{F}}_{0}{\mathbb{L}}_{n}$. Using the Product Formula to $\sigma(\beta^{\prime})^{p^{\lambda}}-\beta^{\prime p^{\lambda}}$, which is $\neq 0$ by (3.8), as in the proof of [1, Lemma 2.1] we get: